OpenStudy (anonymous):

what is the natural numbers between 10 and 40 divisible by 3

OpenStudy (anonymous):

\[ 10\leq 3n \leq 40 \\ 12\leq 3n \leq 39 \\ 4\leq n \leq 13 \]So basically \[\{3n |n\in [4,13]\}\]